Solve using the elimination method. Show your work. If the system has no solution or an infinite number solution state this.
8x + 10y = 92
24x + 11y + 162

Answer by stanbon(75887)

Solve using the elimination method. Show your work. If the system has no solution or an infinite number solution state this.
8x + 10y = 92
24x + 11y = 162
-------------------------------
Multiply thru the 1st equation by 3 to get:
24x + 30y = 3*92
24x + 11y = 162
----------
Subract and solve for "y":
19y = 114
y = 6
-----
Solve for "x":
8x + 10y = 92
8x + 10*6 = 92
8x = 32
x = 4
=================
Cheers,
Stan H.
==================